# FedBaF: Federated Learning Aggregation Biased by a Foundation Model

### v2

@@ 1 Introduction
Developing foundation models (Zhuang et al.,, 2023) has become a major focus for leading technology companies like OpenAI, Microsoft, and Amazon AWS. These deep learning models are often trained with vast amounts of high-quality data (Bommasani et al.,, 2021) and their ability to generalize across different tasks and domains has made them essential assets for industry, government, and academia. Foundation models have been applied to natural language processing (e.g., text generation, translation, summarization), image generation and recognition, healthcare diagnostics, finance predictive analytics, and customer service and virtual assistant tasks (Chen et al.,, 2023; Han et al.,, 2021; Duan et al.,, 2021; Joshi et al.,, 2022; Yosinski et al.,, 2014). When a foundation model’s training data distribution overlaps with a new application, it provides a robust starting point for fine-tuning and customization. Instead of training a model from scratch, with limited data and computes, we can leverage pre-trained foundation models to enable faster training.
For many applications, data that could be used to customize or fine-tune foundation models is often distributed across multiple clients, such as a network of clinics or small companies spread across different jurisdictions. For example, fine-tuning a recommendation model to fit a small company’s product offering may require data from clients in various regions; and adapting a healthcare model for a network of clinics would involve confidential, distributed data sources. Therefore, effective generalization requires access to diverse data from multiple clients (Pranav and Moura,, 2024).
Federated Learning (FL) is a promising solution for fine-tuning these models without sharing client data: FL clients train models on diverse local data, and a central FL server aggregates the client updates to build and refine a global model (Li et al.,, 2021; McMahan et al.,, 2017; Singh et al.,, 2019; Lyu et al.,, 2020; Nguyen et al.,, 2021; Wang et al.,, 2020; Lee et al.,, 2023; Siew et al.,, 2024). Using a foundation model to initialize the global FL model leads to effective customization that leverages diverse, distributed data without directly accessing the client data (Nguyen et al.,, 2022; Chen et al.,, 2023). However, there are significant risks associated with sending a foundation model to clients, as required in traditional FL fine-tuning methods.
@@
First, disclosing a foundation model’s weights to FL clients poses a significant security risk. For example, malicious actors could carry out membership inference attacks, identifying whether specific data was part of the foundation model’s training dataset. Then, an attacker could disrupt the global model’s training by introducing updates that degrade performance on identified data through backdoor attacks, deliberately leading to targeted misclassification (Dayal et al.,, 2023; Hu et al.,, 2022; Wang et al., 2023b, ). Similarly, for model inversion attacks, attackers use known model weights to reverse-engineer sensitive training data (see Figure 1(Fredrikson et al.,, 2015; Li et al., 2022a, ; Zhang et al.,, 2020). Protecting foundation models, often trained on sensitive, proprietary data, is critical for safeguarding the training data and maintaining model integrity (Bagdasaryan et al.,, 2020; Kim et al.,, 2023).
Second, in competitive business contexts, disclosing foundation model weights to clients risks leaking strategic insights and proprietary information to adversaries (Han et al.,, 2021; Yu et al.,, 2023). This undermines a company’s competitive advantage and substantial investments in data collection and training.
@@
When client data distributions are biased or non-IID, e.g., different ratios or a lack of certain labels, clients optimize correspondingly diverse local objective functions (Pranav and Moura,, 2023) and may send conflicting updates to the server that skew the global model (Zhao et al.,, 2018). In FedBaF, the foundation model continuously serves as a form of regularization and stabilizes the global model by reducing the influence of these conflicting updates during the aggregation phase (Chen et al.,, 2023; Li et al., 2022b, ; Tan et al.,, 2022; Yosinski et al.,, 2014).
Furthermore, FedBaF uses a fixed foundation model as an anchor and continuously incorporates it throughout the FL training process. This adds a layer of protection from adversarial attacks beyond those introduced by disclosing foundation model weights – such as misclassification attacks or backdoor attacks, where compromised clients feed malicious updates to the server (Lyu et al.,, 2020; Bagdasaryan et al.,, 2020).
@@
3) We conduct extensive empirical evaluation and show that FedBaF matches or exceeds the training performance of traditional weight initialization methods – with better test performance in 10 out of 14 cases. Our experiments use Pre-ResNet and more complex architectures like Vision Transformer and Transformer-based language models frequently used as foundation models (Xu et al.,, 2023; Kenneweg et al.,, 2024). Compared to standard FedAvg (McMahan et al.,, 2017) and FedProx (Li et al.,, 2020) with foundation models used for weight initialization, FedBaF achieves accuracy improvements of up to 10.8% in IID and up to 37.5% globally and 5.9% locally in non-IID settings. Simultaneously, FedBaF safeguards the foundation model. Similarly, applying FedBaF to a Transformer-based language model significantly reduced perplexity by up to 76.0%.
@@ 2 Related Work
Traditionally, pre-trained models are used in FL to initialize the weights of the global FL model. The server distributes this model to local clients, and the clients update it by using their local data. We refer to this approach as “weight initialization” throughout this paper. Such fine-tuning of a pre-trained model can significantly improve performance of the learned FL global model by integrating data from new clients (Nguyen et al.,, 2022). Several recent studies devised methods that leverage weight initialization to further improve performance: Federated Nearest Class Means (FedNCM) (Legate et al.,, 2023) for last-layer guidance, Federated Recursive Ridge Regression (Fed3R) (Fanì et al.,, 2024), Fractal Pair Similarity (FPS) (Chen et al.,, 2023), and FedPCL (Tan et al.,, 2022).
Several works also explore the use of foundation models in FL. These include cases where a subset of the weights of a large foundation model are chosen to initialize and fine-tune a smaller model (Xu et al.,, 2024) and when clients have diverse model architectures (Wang et al., 2023a, ; Park and Joe-Wong,, 2024). Particularly in scenarios with limited pre-training data (Chen et al.,, 2023), approaches often rely on synthetic data (Nikolenko,, 2021; Chen et al.,, 2023) for pre-training. Federated Prototype-wise Contrastive Learning (FedPCL) is a significant development that improves communication efficiency in FL by using class prototypes (Tan et al.,, 2022) and enhances personalized learning by having clients share class-specific information more effectively.
The related works discussed in this section so far achieve good performance, but they do not consider the significant security vulnerabilities that result from sharing a foundation model with local clients, which compromise data privacy and the integrity of the global model. Malicious clients with access to foundation models can exploit them through: Model Inversion Attacks, recovering original training data or sensitive attributes from the model’s outputs (Fredrikson et al.,, 2015; Zhang et al.,, 2020; Li et al., 2022a, ); Membership Inference Attacks, analyzing model predictions to determine whether specific data records were used in training (Hu et al.,, 2022; Dayal et al.,, 2023; Wang et al., 2023b, ). These attacks compromise the security of the FL system, necessitating the development of more secure methods for leveraging pre-trained foundation models in FL settings.
@@ 3 Methodology
8:     $\mathbf{w}^{\prime}_{t+1}\leftarrow\sum_{k\in S_{t}}\frac{n_{k}}{\sum_{k\in S_{t}}n_{k}}\mathbf{w}_{t+1}^{k}$
@@
Algorithm 1 describes how FedBaF fits into the traditional FL framework by incorporating foundation model weights during aggregation, as illustrated in Lines 9-11 (McMahan et al.,, 2017). FedBaF is versatile and can be embedded into many existing FL algorithms, e.g., SCAFFOLD (Karimireddy et al.,, 2020), FedProx (Li et al.,, 2020), FedAdam (Reddi et al.,, 2021), or other FL strategies, by modifying their aggregation methods (Lines 8-11) and using their existing ClientUpdate$(\mathbf{w},D)$ logic for clients’ local training in Line 13.
@@
Foundation and Global Model Architecture Mismatch. In practice, the foundation model and the global model architectures may differ beyond the input and output layers. These differences can arise in terms of the number of layers or the number of parameters per layer. Since foundation models are typically larger than global FL models – consistent with their role as highly expressive networks pre-trained on extensive data (Meng et al.,, 2023; Awais et al.,, 2025) – we consider the following two cases:
$\bullet$ Foundation model with more layers: When the foundation model is deeper than the global FL model, only a subset of its layers is used during aggregation. We take advantage of the fact that most large models have layers grouped into sections. Here, a section is a contiguous subset of layers within a model that shares similar structural properties, such as the number of parameters, functional roles, or connectivity patterns. FedBaF selects and matches sections between the foundation and global models, prioritizing feature extraction sections near the input layer. Within each matched section, only layers that align with the global model’s architecture are integrated into FL training. Methods for selecting compatible layers during aggregation are explored in Park and Joe-Wong, (2024); Xu et al., (2024).
@@
Designing $\tau_{t}$. Our careful design of $\tau_{t}$ uses the $L2$ norm of the difference between consecutive normalized weights of $\mathbf{w^{\prime}}_{t+1}$ and $\mathbf{w}_{t}$, divided by $\sqrt{t+1}$. This change in the model’s weights between rounds reflects how much the global model adapts to new client updates. The normalization prevents $\tau_{t}$ from becoming too large. The factor $\sqrt{t+1}$ ensures that, as training progresses, the influence of $\mathbf{w}_{pre}$ gradually diminishes, but not too quickly, and $\mathbf{w}_{t+1}$ approaches the improving averaged weights $\mathbf{w^{\prime}}_{t+1}$. This strategy is critical to keeping the global model flexible and effective, especially when client data differs from the data used to train the foundation model (Karimireddy et al.,, 2020).
@@ 4 THEORETICAL ANALYSIS
where $\mathbf{w}$ are model (usually, deep neural network) weights and the $f_{\mathcal{D}_{k}}$ are $L$-smooth local objective functions. The convergence analysis presented in this section also makes the following standard assumptions made by (Karimireddy et al.,, 2020) and detailed in Appendix E. Each client locally optimizes $\mathbf{w}$ using stochastic gradient descent, where the stochastic gradients are (i) unbiased and (ii) have bounded variance. We also assume (iii) bounded gradient dissimilarity: the norm of the difference between the gradient of the global objective and the gradients computed using different local objective functions is bounded. Lastly, we assume that (iv) the foundation model has the same architecture as the global model, ensuring compatibility during aggregation. See Appendix E.1.1 for mathematical details regarding the assumptions.
@@ 4.1 General Convergence Analysis
Using the same restrictions on local and global learning rates placed by the FedAvg convergence analysis in (Karimireddy et al.,, 2020), the aforementioned bounded gradient variance, bounded gradient dissimilarity, and L-smoothness assumptions ensure that our method converges to $\mathbf{w}^{*}$ faster than FedAvg. Similar convergence rate arguments for other FL methods with appropriately modified aggregation can be shown, as discussed in Sec. 3.
@@ 5.1 Experimental Setup
Our experiments with popular image classification tasks encompass experiments using the CIFAR-10 and Rome Weather Image (Vaz,, 2021) datasets. We use Pre-ResNet and Vision Transformer (Dosovitskiy et al.,, 2021) architectures as Vision Transformers are popular foundation model architectures known for achieving remarkable performance (Zhou et al.,, 2024) for ImageNet challenges (Dosovitskiy et al.,, 2021). We train the foundation models on the Tiny ImageNet-200 and Weather Image (Xiao,, 2021) datasets. Our evaluations consider both IID and non-IID settings (see Appendix A). To demonstrate FedBaF’s generalizability to other tasks, we also evaluate it on a next-word prediction task using a Transformer language model pre-trained on the WikiText-2 dataset and tested on the Penn Treebank dataset.
@@ 5.2 Experimental Results
Figures 3, 4, and 5 display the extensive test accuracy/perplexity evaluation results for FedBaF and our two baselines (no foundation model and weight initialization). We evaluate the three methods using both FedAvg and FedProx (Li et al.,, 2020) as the base FL training algorithms. We incorporate FedBaF into FedProx by modifying Line 8 and the ClientUpdate routine in Algorithm 1, including both IID and non-IID settings, with one non-adversarial scenario and four adversarial scenarios. We use FedProx’s aggregation step: $\mathbf{w}^{\prime}_{t+1}\leftarrow\sum_{k\in S_{t}}\frac{n_{k}}{\sum_{k\in S_{t}}n_{k}}\mathbf{w}_{t+1}^{k}-\mu(\mathbf{w}_{t+1}^{k}-\mathbf{w}_{t})$. Here, $\mu$ represents the regularization term that controls the trade-off between the local and global objectives. FedBaF then incorporates the foundation model weights as in Lines 9-11 of Algorithm 1.
@@ 5.2.1 Testing FedBaF Performance
In Appendix C.3, we provide additional examples of FedBaF achieving privacy and test performance improvements. They include comparisons with FedAdam (Reddi et al.,, 2021), a state-of-the-art FL method leveraging pre-trained models.
@@ 6 Conclusion
This section includes supplementary experimental results, analyzing the effect of varying foundation model quality on FedBaF’s performance and comparisons using the official pre-trained foundation model. Computational complexity is also compared to scenarios without foundation models and weight initialization. Additionally, we include further experiments employing FedAdam (Reddi et al.,, 2021).
@@ C.3 Generalized Applicability of FedBaF with FedAdam
To further assess the generalized applicability of FedBaF, we conducted additional experiments using the FedAdam Reddi et al., (2021) algorithm. Specifically, we evaluated its performance on image classification tasks with CIFAR-10 and Rome Weather Image datasets using Pre-ResNet and Vision Transformer, as well as on a language modeling task with a Transformer model pre-trained on the WikiText-2 dataset. We followed the same experimental setup, attack scenarios, and hyperparameters described in Section 5 and Appendix A.

### v3

@@ 1 Introduction
Developing foundation models (ZCL, 23) has become a major focus for leading technology companies like OpenAI, Microsoft, and Amazon AWS. These deep learning models are often trained with vast amounts of high-quality data (BHA+, 21) and their ability to generalize across different tasks and domains has made them essential assets for industry, government, and academia. Foundation models have been applied to natural language processing (e.g., text generation, translation, summarization), image generation and recognition, healthcare diagnostics, finance predictive analytics, and customer service and virtual assistant tasks (CTL+, 23; HZD+, 21; DLJ+, 21; JPS, 22; YCBL, 14). When a foundation model’s training data distribution overlaps with a new application, it provides a robust starting point for fine-tuning and customization. Instead of training a model from scratch, with limited data and computes, we can leverage pre-trained foundation models to enable faster training.
For many applications, data that could be used to customize or fine-tune foundation models is often distributed across multiple clients, such as a network of clinics or small companies spread across different jurisdictions. For example, fine-tuning a recommendation model to fit a small company’s product offering may require data from clients in various regions; and adapting a healthcare model for a network of clinics would involve confidential, distributed data sources. Therefore, effective generalization requires access to diverse data from multiple clients (PM, 24).
Federated Learning (FL) is a promising solution for fine-tuning these models without sharing client data: FL clients train models on diverse local data, and a central FL server aggregates the client updates to build and refine a global model (LWW+, 21; MMR+, 17; SVGR, 19; LYZY, 20; NDP+, 21; WYS+, 20; LZA, 23; SZP+, 24). Using a foundation model to initialize the global FL model leads to effective customization that leverages diverse, distributed data without directly accessing the client data (NMSR, 22; CTL+, 23). However, there are significant risks associated with sending a foundation model to clients, as required in traditional FL fine-tuning methods.
@@
First, disclosing a foundation model’s weights to FL clients poses a significant security risk. For example, malicious actors could carry out membership inference attacks, identifying whether specific data was part of the foundation model’s training dataset. Then, an attacker could disrupt the global model’s training by introducing updates that degrade performance on identified data through backdoor attacks, deliberately leading to targeted misclassification (DAATM, 23; HSS+, 22; WWW+, 23). Similarly, for model inversion attacks, attackers use known model weights to reverse-engineer sensitive training data (see Figure 1(FJR, 15; LRC+, 22; ZJP+, 20). Protecting foundation models, often trained on sensitive, proprietary data, is critical for safeguarding the training data and maintaining model integrity (BVH+, 20; KLM+, 23).
Second, in competitive business contexts, disclosing foundation model weights to clients risks leaking strategic insights and proprietary information to adversaries (HZD+, 21; YWZ+, 23). This undermines a company’s competitive advantage and substantial investments in data collection and training.
@@
When client data distributions are biased or non-IID, e.g., different ratios or a lack of certain labels, clients optimize correspondingly diverse local objective functions (PM, 23) and may send conflicting updates to the server that skew the global model (ZLL+, 18). In FedBaF, the foundation model continuously serves as a form of regularization and stabilizes the global model by reducing the influence of these conflicting updates during the aggregation phase (CTL+, 23; LRJ+, 22; TLM+, 22; YCBL, 14).
Furthermore, FedBaF uses a fixed foundation model as an anchor and continuously incorporates it throughout the FL training process. This adds a layer of protection from adversarial attacks beyond those introduced by disclosing foundation model weights – such as misclassification attacks or backdoor attacks, where compromised clients feed malicious updates to the server (LYZY, 20; BVH+, 20).
@@
3) We conduct extensive empirical evaluation and show that FedBaF matches or exceeds the training performance of traditional weight initialization methods – with better test performance in 10 out of 14 cases. Our experiments use Pre-ResNet and more complex architectures like Vision Transformer and Transformer-based language models frequently used as foundation models (XZC, 23; KKH, 24). Compared to standard FedAvg (MMR+, 17) and FedProx (LSZ+, 20) with foundation models used for weight initialization, FedBaF achieves accuracy improvements of up to 10.8% in IID and up to 37.5% globally and 5.9% locally in non-IID settings. Simultaneously, FedBaF safeguards the foundation model. Similarly, applying FedBaF to a Transformer-based language model significantly reduced perplexity by up to 76.0%.
@@ 2 Related Work
Traditionally, pre-trained models are used in FL to initialize the weights of the global FL model. The server distributes this model to local clients, and the clients update it by using their local data. We refer to this approach as “weight initialization” throughout this paper. Such fine-tuning of a pre-trained model can significantly improve performance of the learned FL global model by integrating data from new clients (NMSR, 22). Several recent studies devised methods that leverage weight initialization to further improve performance: Federated Nearest Class Means (FedNCM) (LBPC+, 23) for last-layer guidance, Federated Recursive Ridge Regression (Fed3R) (FCCC, 24), Fractal Pair Similarity (FPS) (CTL+, 23), and FedPCL (TLM+, 22).
Several works also explore the use of foundation models in FL. These include cases where a subset of the weights of a large foundation model are chosen to initialize and fine-tune a smaller model (XCV+, 24) and when clients have diverse model architectures (WHC+, 23; PJW, 24). Particularly in scenarios with limited pre-training data (CTL+, 23), approaches often rely on synthetic data (Nik, 21; CTL+, 23) for pre-training. Federated Prototype-wise Contrastive Learning (FedPCL) is a significant development that improves communication efficiency in FL by using class prototypes (TLM+, 22) and enhances personalized learning by having clients share class-specific information more effectively.
The related works discussed in this section so far achieve good performance, but they do not consider the significant security vulnerabilities that result from sharing a foundation model with local clients, which compromise data privacy and the integrity of the global model. Malicious clients with access to foundation models can exploit them through: Model Inversion Attacks, recovering original training data or sensitive attributes from the model’s outputs (FJR, 15; ZJP+, 20; LRC+, 22); Membership Inference Attacks, analyzing model predictions to determine whether specific data records were used in training (HSS+, 22; DAATM, 23; WWW+, 23). These attacks compromise the security of the FL system, necessitating the development of more secure methods for leveraging pre-trained foundation models in FL settings.
@@ 3 Methodology
8:     $\mathbf{w}^{\prime}_{t+1}\leftarrow\sum_{k\in S_{t}}\frac{n_{k}}{\sum_{\ell\in S_{t}}n_{\ell}}\mathbf{w}_{t+1}^{k}$
@@
Algorithm 1 describes how FedBaF fits into the traditional FL framework by incorporating foundation model weights during aggregation, as illustrated in Lines 9-11 (MMR+, 17). FedBaF is versatile and can be embedded into many existing FL algorithms, e.g., SCAFFOLD (KKM+, 20), FedProx (LSZ+, 20), FedAdam (RCZ+, 21), or other FL strategies, by modifying their aggregation methods (Lines 8-11) and using their existing ClientUpdate$(\mathbf{w},D)$ logic for clients’ local training in Line 13.
@@
Foundation and Global Model Architecture Mismatch. In practice, the foundation model and the global model architectures may differ beyond the input and output layers. These differences can arise in terms of the number of layers or the number of parameters per layer. Since foundation models are typically larger than global FL models – consistent with their role as highly expressive networks pre-trained on extensive data (MSJ+, 23; ANK+, 25) – we consider the following two cases:
$\bullet$ Foundation model with more layers: When the foundation model is deeper than the global FL model, only a subset of its layers is used during aggregation. We take advantage of the fact that most large models have layers grouped into sections. Here, a section is a contiguous subset of layers within a model that shares similar structural properties, such as the number of parameters, functional roles, or connectivity patterns. FedBaF selects and matches sections between the foundation and global models, prioritizing feature extraction sections near the input layer. Within each matched section, only layers that align with the global model’s architecture are integrated into FL training. Methods for selecting compatible layers during aggregation are explored in PJW (24); XCV+ (24).
@@
Designing $\tau_{t}$. Our careful design of $\tau_{t}$ uses the $L2$ norm of the difference between consecutive normalized weights of $\mathbf{w^{\prime}}_{t+1}$ and $\mathbf{w}_{t}$, divided by $\sqrt{t+1}$. This change in the model’s weights between rounds reflects how much the global model adapts to new client updates. The normalization prevents $\tau_{t}$ from becoming too large. The factor $\sqrt{t+1}$ ensures that, as training progresses, the influence of $\mathbf{w}_{pre}$ gradually diminishes, but not too quickly, and $\mathbf{w}_{t+1}$ approaches the improving averaged weights $\mathbf{w^{\prime}}_{t+1}$. This strategy is critical to keeping the global model flexible and effective, especially when client data differs from the data used to train the foundation model (KKM+, 20).
@@ 4 THEORETICAL ANALYSIS
where $\mathbf{w}$ are model (usually, deep neural network) weights and the $f_{\mathcal{D}_{k}}$ are $L$-smooth local objective functions. The convergence analysis presented in this section also makes the following standard assumptions made by (KKM+, 20) and detailed in Appendix E. Each client locally optimizes $\mathbf{w}$ using stochastic gradient descent, where the stochastic gradients are (i) unbiased and (ii) have bounded variance. We also assume (iii) bounded gradient dissimilarity: the norm of the difference between the gradient of the global objective and the gradients computed using different local objective functions is bounded. Lastly, we assume that (iv) the foundation model has the same architecture as the global model, ensuring compatibility during aggregation. See Appendix E.1.1 for mathematical details regarding the assumptions.
@@ 4.1 General Convergence Analysis
Using the same restrictions on local and global learning rates placed by the FedAvg convergence analysis in (KKM+, 20), the aforementioned bounded gradient variance, bounded gradient dissimilarity, and L-smoothness assumptions ensure that our method converges to $\mathbf{w}^{*}$ faster than FedAvg. Similar convergence rate arguments for other FL methods with appropriately modified aggregation can be shown, as discussed in Sec. 3.
@@ 5.1 Experimental Setup
Our experiments with popular image classification tasks encompass experiments using the CIFAR-10 and Rome Weather Image (Vaz, 21) datasets. We use Pre-ResNet and Vision Transformer (DBK+, 21) architectures as Vision Transformers are popular foundation model architectures known for achieving remarkable performance (ZLL+, 24) for ImageNet challenges (DBK+, 21). We train the foundation models on the Tiny ImageNet-200 and Weather Image (Xia, 21) datasets. Our evaluations consider both IID and non-IID settings (see Appendix A). To demonstrate FedBaF’s generalizability to other tasks, we also evaluate it on a next-word prediction task using a Transformer language model pre-trained on the WikiText-2 dataset and tested on the Penn Treebank dataset.
@@ 5.2 Experimental Results
Figures 3, 4, and 5 display the extensive test accuracy/perplexity evaluation results for FedBaF and our two baselines (no foundation model and weight initialization). We evaluate the three methods using both FedAvg and FedProx (LSZ+, 20) as the base FL training algorithms. We incorporate FedBaF into FedProx by modifying Line 8 and the ClientUpdate routine in Algorithm 1, including both IID and non-IID settings, with one non-adversarial scenario and four adversarial scenarios. We use FedProx’s aggregation step: $\mathbf{w}^{\prime}_{t+1}\leftarrow\sum_{k\in S_{t}}\frac{n_{k}}{\sum_{k\in S_{t}}n_{k}}\mathbf{w}_{t+1}^{k}-\mu(\mathbf{w}_{t+1}^{k}-\mathbf{w}_{t})$. Here, $\mu$ represents the regularization term that controls the trade-off between the local and global objectives. FedBaF then incorporates the foundation model weights as in Lines 9-11 of Algorithm 1.
@@ 5.2.1 Testing FedBaF Performance
In Appendix C.3, we provide additional examples of FedBaF achieving privacy and test performance improvements. They include comparisons with FedAdam (RCZ+, 21), a state-of-the-art FL method leveraging pre-trained models.
@@ 6 Conclusion
This section includes supplementary experimental results, analyzing the effect of varying foundation model quality on FedBaF’s performance and comparisons using the official pre-trained foundation model. Computational complexity is also compared to scenarios without foundation models and weight initialization. Additionally, we include further experiments employing FedAdam [31].
@@ C.3 Generalized Applicability of FedBaF with FedAdam
To further assess the generalized applicability of FedBaF, we conducted additional experiments using the FedAdam [31] algorithm. Specifically, we evaluated its performance on image classification tasks with CIFAR-10 and Rome Weather Image datasets using Pre-ResNet and Vision Transformer, as well as on a language modeling task with a Transformer model pre-trained on the WikiText-2 dataset. We followed the same experimental setup, attack scenarios, and hyperparameters described in Section 5 and Appendix A.